# LLM4AD: Large Language Models for Autonomous Driving — Concept, Review, Benchmark, Experiments, and Future Trends

### v4

@@ V-D2 The Improvement on Personalization Performance
This section explores the impact of integrating the Talk2Drive framework into autonomous driving systems to enhance personalization based on the LLM model GPT-4 [openai_gpt-4_2023]. One of the central focuses of our system is its ability to offer drivers a personalized driving experience. Through Talk2Drive, individuals can express their preferences or feelings with varying degrees of directness, prompting the system to make corresponding adjustments. The commands collected are divided into three levels of directness, as explained in Tab. VII.
For each command directness within each expected driving behavior, the system’s responses to various responses given by human participants are collected and calculated into the defined metrics. Here, we utilize the takeover rate as our primary metric and evaluate it both pre- and post-integration of Talk2Drive. We incorporate a rule-based system with a keyword-trigger logic as the baseline for comparison. For instance, in the highway scenario, commands like “accelerate,” “left,” and “right” prompt the system to execute relevant actions such as speeding up or changing lanes accordingly. The results are gathered in diverse driving scenarios on human drivers with diverse driving styles.
@@
As shown in Tab. IX, the takeover rate in all driving scenarios decreases significantly, ranging from a 56.3% reduction to the complete elimination of takeover cases, demonstrating the Talk2Drive system’s ability to personalize the driving experience based on human preference. Additionally, one important finding from our results is that for all levels of command directness, our Talk2Drive system shows significant improvements compared to the baseline systems.
@@ V-D3 The Effectiveness of the Memory Module
To further investigate the performance of our memory module, we also conduct experiments comparing personalization across two settings: conditions without the memory module, and conditions with the memory module. These experiments take place in the parking lot, where the vehicle follows a pre-defined trajectory and adjusts speed based on human inputs. Humans take over in this context when they find the adjusted speed unsatisfactory. As demonstrated in Tab. X, there are significant reductions in takeover rate with Talk2Drive, regardless of the driver’s aggression or conservatism levels.
@@
Driver
Without Memory Module
With Memory Module
@@ VI Real-Vehicle Experiments: On-Board VLM for Personalized Motion Control
Figure 11: The proposed framework for personalized motion control with an on-board vision language model in real-vehicle experiments.
@@ VI-C Real-World Experiment
Lane
Change
Baseline 2.44 3.91 1.65 0.37 3.14 0.88 1.01 - 88.5 60.0 74.5
@@
Left
Turn
Baseline - 1.12 7.52 0.22 1.81 1.29 1.36 - 88.0 60.0 70.4
@@
Figure 12: Overview of the experiment track.
@@ VI-C4 Ablation Study on Memory Module
Figure 15: Effectiveness of memory modules in takeover rates
@@ IX Conclusions
[Uncaptioned image] Can Cui (Student Member, IEEE) received his M.S. degree in Electrical and Computer Engineering from the University of Michigan, Ann Arbor in 2022. He is now pursuing a Ph.D. degree in the College of Engineering at Purdue University. His research pursuits encompass automated driving, Advanced Driver-Assistance Systems (ADAS), applied machine learning, controls, and digital twin.
[Uncaptioned image] Yunsheng Ma (Student Member, IEEE) received his M.S. degree in Computer Science from New York University in 2022. He is now pursuing a Ph.D. degree in the College of Engineering at Purdue University, with a research focus on autonomous driving, embodied AI, and digital twins. He has publications in top-tier conferences and journals, including CVPR, AAAI, ITSC, and IEEE Transactions on Intelligent Vehicles.
[Uncaptioned image] Sung-Yeon Park received her M.S. degree in Civil Engineering from Seoul National University in 2023. She is now pursuing a Ph.D. degree in the College of Engineering at Purdue University, with a research focus on autonomous driving for 3D/4D scene understanding and spatial reasoning.
[Uncaptioned image] Zichong Yang received his M.S. degree in Mechanical Engineering from Carnegie Mellon University in 2023. He is now pursuing a Ph.D. degree in the College of Engineering at Purdue University, with a research focus on computer vision, human pose estimation, and human trajectory prediction.
[Uncaptioned image] Yupeng Zhou (Student Member, IEEE) received his M.S. degree in Mechanical Engineering from Purdue University, West Lafayette, in 2024. He is now a Ph.D. student in the College of Engineering at Purdue University, with a research focus on applied robotics, autonomous vehicles, and controls.
[Uncaptioned image] Juanwu Lu (Student Member, IEEE) received an M.S. degree in Civil Engineering from the University of California, Berkeley in 2022. He is currently pursuing his Ph.D. degree in the College of Engineering at Purdue University. His research interests include statistical machine learning and its application for generalizable and interpretable microscopic traffic modeling.
[Uncaptioned image] Juntong Peng (Student Member, IEEE) received his B.Eng. degree in Information Engineering from Shanghai Jiao Tong University in 2024. He is now pursuing his Ph.D. degree in the College of Engineering at Purdue University, with a research focus on autonomous driving, collaborative perception, and multi-agent intelligence.
[Uncaptioned image] Jiaru Zhang received his Ph.D. degree in Computer Science from Shanghai Jiao Tong University in 2024. He is currently a postdoctral scholar in the Institute for Physical Artificial Intelligence at Purdue University. His research interests include Bayesian deep learning, generative models, and their applications in autonomous driving.
[Uncaptioned image] Ruqi Zhang received her Ph.D. degree from Cornell University. She is currently an assistant professor in the Department of Computer Science at Purdue University. Before that, she was a postdoctoral fellow at the Institute for Foundations of Machine Learning, UT Austin. Her research focuses on machine learning, deep learning, generative modeling, probabilistic methods, and trustworthy AI.
[Uncaptioned image] Lingxi Li (Senior Member, IEEE) received his Ph.D. degree in Electrical and Computer Engineering from the University of Illinois at Urbana–Champaign in 2008. He is a professor in Electrical and Computer Engineering at Purdue University. His research interests include complex systems, intelligent transportation systems, connected and automated vehicles, discrete event dynamic systems, human-machine interaction, digital twins, and parallel intelligence.
[Uncaptioned image] Yaobin Chen (Senior Member, IEEE) received his Ph.D. degree in Electrical and Computer Engineering from Rensselaer Polytechnic Institute in 1988. He is a professor in Electrical and Computer Engineering at Purdue University. His areas of expertise include automated driving and autonomous/connected vehicles, vehicle active safety systems, EV and HEV, intelligent controls and robotics, and advanced energy and power systems.
[Uncaptioned image] Jitesh Panchal received his Ph.D. degree in Mechanical Engineering from Georgia Institute of Technology in 2005. He is a professor in Mechanical Engineering at Purdue University. His areas of expertise include computational design of socio-technical systems, secure design and manufacturing, engineering design by self-organized virtual communities, and integrated products and materials design.
[Uncaptioned image] Amr Abdelraouf (Member, IEEE) received his Ph.D. degree in Transportation Engineering from the University of Central Florida in 2022. He was a research scientist at Toyota Motor North America, InfoTech Labs. His research interests include applied AI/ML, advanced driver assistance systems, connected and automated vehicles, intelligent transportation systems, and digital twins.
[Uncaptioned image] Rohit Gupta received the Ph.D. degree in Mechanical Engineering from the University of California at Santa Barbara. He is currently a principal researcher at Toyota Motor North America, InfoTech Labs. Prior to joining Toyota, he worked in various industries: semiconductors, telecommunications, finance, and software. His research interests include cloud and edge computing architectures, connected and autonomous vehicles, AI/ML, “Digital Twin” of human-vehicle interactions, and high-speed mobility.
[Uncaptioned image] Kyungtae (KT) Han (Senior Member, IEEE) received the Ph.D. degree in Electrical and Computer Engineering from The University of Texas at Austin in 2006. He is currently a senior principal scientist at Toyota Motor North America, InfoTech Labs. Prior to joining Toyota, Dr. Han was a Research Scientist at Intel Labs and a Director at Locix Inc. His research interests include cyber-physical systems, connected and automated vehicle techniques, and intelligent transportation systems.
[Uncaptioned image] Ziran Wang (Senior Member, IEEE) received his Ph.D. degree in Mechanical Engineering from the University of California, Riverside in 2019. He is an assistant professor in the College of Engineering at Purdue University, and was a principal researcher at Toyota Motor North America R&D. His research focuses on automated driving, human-autonomy teaming, and digital twins. He is the Founding Chair of the IEEE Technical Committee on Internet of Things in Intelligent Transportation Systems.

### v5

@@ V-D2 The Improvement on Personalization Performance
This section explores the impact of integrating the Talk2Drive framework into autonomous driving systems to enhance personalization based on the LLM model GPT-4 [openai_gpt-4_2023]. One of the central focuses of the system is its ability to offer drivers a personalized driving experience. Through Talk2Drive, individuals can express their preferences or feelings with varying degrees of directness, prompting the system to make corresponding adjustments. The commands collected are divided into three levels of directness, as explained in Tab. VII.
For each command directness within each expected driving behavior, the system’s responses to various responses given by human participants are collected and calculated into the defined metrics. Here, the takeover rate is utilized as the primary metric and is evaluated both pre- and post-integration of Talk2Drive. We incorporate a rule-based system with a keyword-trigger logic as the baseline for comparison. For instance, in the highway scenario, commands like “accelerate,” “left,” and “right” prompt the system to execute relevant actions such as speeding up or changing lanes accordingly. The results are gathered in diverse driving scenarios on human drivers with diverse driving styles.
@@
As shown in Tab. IX, the takeover rate in all driving scenarios decreases significantly, ranging from a 56.3% reduction to the complete elimination of takeover cases, demonstrating the Talk2Drive system’s ability to personalize the driving experience based on human preference. Additionally, one important finding from the results is that for all levels of command directness, the Talk2Drive system shows significant improvements compared to the baseline systems.
@@ V-D3 The Effectiveness of the Memory Module
To further investigate the performance of the memory module, experiments comparing personalization across two settings: conditions without the memory module, and conditions with the memory module are conducted. These experiments take place in the parking lot, where the vehicle follows a pre-defined trajectory and adjusts speed based on human inputs. Humans take over in this context when they find the adjusted speed unsatisfactory. As demonstrated in Tab. X, there are significant reductions in takeover rate with Talk2Drive, regardless of the driver’s aggression or conservatism levels.
@@
Driver Without Memory Module With Memory Module
@@ VI Real-Vehicle Experiments: On-Board VLM for Personalized Motion Control
Figure 9: The proposed framework for personalized motion control with an on-board vision language model in real-vehicle experiments. The system processes four primary inputs: system messages, human instructions, camera images, and historical context retrieved from a retrieval-augmented generation (RAG) memory module. Based on these inputs, the model generates an action policy consisting of specific matrices that adjust the parameters for model predictive control (lateral) and PID control (longitudinal). These adjusted parameters regulate steering and acceleration through a drive-by-wire system to execute the driving maneuvers. The process concludes with a feedback loop where human evaluations are committed to the memory module, updating the historical database to refine future personalization.
@@ VI-C Real-World Experiment
Lane Change Baseline 2.44 3.91 1.65 0.37 3.14 0.88 1.01 - 88.5 60.0 74.5
@@
Left Turn Baseline - 1.12 7.52 0.22 1.81 1.29 1.36 - 88.0 60.0 70.4
@@
Figure 10: This figure provides an overview of the closed-loop test track layout used for personalized motion control real-world validation experiments.
@@ VI-C4 Ablation Study on Memory Module
Figure 13: Effectiveness of memory modules in takeover rates.